# Policy Contrastive Decoding for Robotic Foundation Models

### v2

@@ Abstract
Robotic foundation models, or generalist robot policies, hold immense potential to enable flexible, general-purpose and dexterous robotic systems. Despite their advancements, our empirical experiments reveal that existing robot policies are prone to learning spurious correlations from pre-training trajectories, adversely affecting their generalization capabilities beyond the training data. To tackle this, we propose a novel Policy Contrastive Decoding (PCD) approach, which redirects the robot policy’s focus toward object-relevant visual clues by contrasting action probability distributions derived from original and object-masked visual inputs. As a training-free method, our PCD can be used as a plugin to improve different types of robot policies without needing to finetune or access model weights. We conduct extensive experiments on top of three open-source robot policies, including the autoregressive policy OpenVLA and the diffusion-based policies Octo and $\pi_{0}$. The obtained results in both simulation and real-world environments prove PCD’s flexibility and effectiveness, e.g., PCD enhances the state-of-the-art policy $\pi_{0}$ by 8% in the simulation environment and by 108% in the real-world environment.
@@ 1 Introduction
Recent efforts in developing flexible, general-purpose, and dexterous robotic systems have largely focused on robotic foundation models, or generalist robot policies. The overarching goal of such policies is to empower users to instruct robots to perform arbitrary tasks, enabling autonomous task execution with minimal human supervision. Specifically, the policy takes as input a visual observation of the robot’s state combined with a language instruction that defines the task, and outputs a robot action, such as end-effector displacement. Ongoing advancements in scaling real-world robotic data corpora have facilitated the development of numerous robot policies [4, 20, 32, 36], which have demonstrated exceptional effectiveness in controlling various robots across diverse environments and acquiring a wide range of manipulation skills.
Despite their notable achievements, existing robot policies are prone to learning spurious correlations [49, 13, 25] from pre-training trajectories, adversely affecting their ability to generalize outside their training data. As illustrated in Fig. 1, the robot policy predominantly relies on spurious features (e.g., background or textures) rather than object features within observations to predict actions (see (a)(d)). Consequently, slight alterations to the visual observation background, like adjustments to the panning light region (i.e., ($\xrightarrow{}$(b)) and the handle position (i.e., ($\xrightarrow{}$(c)), lead to 36% and 32% drops in the policy’s action prediction performance, respectively. Prior work in other fields has demonstrated that models relying on spurious correlations often suffer significant performance degradation on test data when a distribution shift occurs between the training and test environments [31, 44, 17]. This underscores the importance of redirecting robot policies’ focus from spurious features to object-relevant ones to enhance their generalization across different scenarios during inference. Therefore, we provide a first study on the following question: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] Can we propose a training-free, plug-and-play approach to mitigate the adverse effects of spurious correlations, thereby rectifying predicted actions for existing robot policies?
@@
In this work, we present a Policy Contrastive Decoding (PCD) approach to answer the question. The core idea of our PCD approach is to shift the robot policy’s attention from spurious features to object-relevant ones by contrasting action probability distributions derived from original and object-masked observation images during inference. To ensure both effectiveness and flexibility of the approach, a Tracking-to-Mask (Track2Mask) strategy is introduced, which first annotates target-specific objects in the initial visual observation using human expertise (e.g., Point and Box prompts) or off-the-shelf object detection models (e.g., Grounding DINO [28]), then employs the SAM2 [21] model to track these objects across subsequent observations in the trajectory. This enables precise object masking with minimal or no human intervention. Moreover, unlike autoregressive policies, diffusion-based policies cannot directly generate action probability distributions to feed into our PCD. To cope with this limitation, we introduce a KDE-based Probabilistic Modeling (KDE-PM) scheme, which calculates action probability distributions for diffusion-based policies using the kernel density estimation (KDE) [47] technique. This enables our PCD approach to be compatible with both autoregressive and diffusion-based policies.
@@
Figure 1: Robot policies tend to spuriously correlate task-irrelevant features with actions, compromising their ability to generalize to unseen scenarios. As observed, changing the light position from (a) to (b) and the drawer handle position from (a) to (c) results in 36% and 32% drops in the performance of the baseline policy OpenVLA [20], respectively. (d) Attention map. More results are in Section 4.4 and Appendix A.1.
@@
To the best of our knowledge, this is the first work to introduce a training-free approach for resolving the spurious correlation issue in robot policies. We conduct extensive evaluations in both simulation and real-world environments, across a total of 15 diverse tasks, and on top of three open-source robot policies—including the autoregressive policy OpenVLA [20] and the diffusion-based policies Octo [32] and $\pi_{0}$ [3]. The achieved results demonstrate the strong flexibility and effectiveness of our PCD approach. On simulation benchmarks, PCD enhances the performance of Octo, OpenVLA and $\pi_{0}$ by 31%, 41% and 8%, respectively. On real-world manipulation tasks, PCD outperforms the state-of-the-art robot policy $\pi_{0}$ by 108%.
@@
Contributions. Our contributions in this work are threefold. 1) We propose PCD, a simple, training-free, and easy-to-implement scheme to tackle the spurious correlation issue in robot policies. 2) PCD can be used as a plugin to enhance both autoregressive and diffusion-based policies, without requiring fine-tuning or access to pre-trained model weights. 3) Extensive experiments demonstrate PCD’s effectiveness and flexibility, consistently improving performance over three state-of-the-art policies across 15 manipulation tasks on multiple robotic platforms.
@@ 2 Related Work
Robotic Foundation Models. The rapid development of robotic foundation models or generalist robot policies has been significantly inspired by the success of large visual and language models [18, 52, 42]. Robot policies can be categorized based on action prediction strategies into autoregressive models and diffusion-based policies. Autoregressive policies sequentially generate action tokens based on previous outputs [4, 5, 9, 50, 1, 53, 54]. OpenVLA [20] equips with a Llama 2 [42] language model alongside a visual encoder that merges pretrained features from SigLIP [52] and DINO-v2 [34]. Diffusion-based policies simultaneously generate entire action dimensions by performing diffusion processes on noise [32, 27]. Octo [40] learns a lightweight policy on the OXE dataset [9], enabling rapid adaptation to unknown tasks and scenarios on standard consumer-grade GPUs. $\pi_{0}$ [3] exhibits flexible control over different robot types by leveraging pre-trained general knowledge from vision-language models. Although most robot policies are trained on extensive robotic behavior datasets, recent studies have shown that slight variations in deployment scenarios can significantly affect their generalization performance on downstream tasks in real-world environments [11, 12, 29].
@@
Contrastive Decoding. Recent studies in large vision-language models (VLMs) have highlighted the challenge of hallucination [14, 37, 26], where models generate inaccurate or misleading outputs not grounded in the input data. Contrastive Decoding (CD) [10, 7, 22] suppresses hallucinations by comparing output distributions from original and distorted input without altering the model’s weights. Instruction Contrastive Decoding (ICD) [46] is used to combat hallucinations in multimodal tasks by modulating the confidence in multimodal alignment of the model’s visual and textual inputs, enabling it to distinguish between hallucinated and relevant tokens. Visual Contrastive Decoding (VCD) [35, 10] aims to reduce object hallucinations by comparing outputs from original and distorted visual inputs. This approach provides a computationally efficient solution as it requires neither additional training nor external pre-trained models. Our proposed Policy Contrastive Decoding (PCD) approach in this work is mainly inspired by the success of VCD schemes. However, unlike those schemes that contrast output distributions from original and noise-distorted [35] or image-removed [10] inputs to counteract LVLMs’ overreliance on language priors, our PCD contrasts action probability distributions derived from original and object-masked inputs, thereby redirecting the robot policy’s focus to object-relevant visual features and enhancing their generalization across different scenarios.
@@ 3.1 Preliminaries
This work focuses on robot manipulation tasks where the objective is to leverage a learned language-conditioned policy $\pi_{\theta}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i},\ell)$ to predict an $M$-dimensional action $\boldsymbol{a}_{i}=[a_{1},...,a_{M}]\in\mathbb{R}^{M}$ conditioned on the current visual observation $\boldsymbol{o}_{i}$ and the language instruction $\ell$ of the task. In this part, we formalize two mainstream paradigms for action prediction: autoregressive action prediction (e.g., OpenVLA [20]) and diffusion-based action prediction (e.g., Octo [32] and $\pi_{0}$ [3]).
@@
Figure 2: Overview of our Proposed Policy Contrastive Decoding (PCD) approach. PCD serves as a plugin to redirect the robot policy’s focus toward object-relevant visual cues by contrasting action probability distributions derived from original observations ${p}$ and object-masked observations $\hat{p}$. For illustrative purposes, we visualize the predictions only in the $\Delta x$ and $\Delta y$ dimensions of the robot action space [$\Delta x$, $\Delta y$, $\Delta z$, rotx, roty, rotz, gripper].
@@
Autoregressive Robot Policies. The policy decomposes the action prediction task into sequential token generation. Each action dimension $a_{t}$ of $\boldsymbol{a}_{i}$ is sampled autoregressively from the probability distribution, conditioned on the visual observation $\boldsymbol{o}_{i}$, the language instruction $\ell$, as well as the previously predicted dimensions $a_{<t}=\{{a}_{t}\}_{t=1}^{t-1}$:
$a_{t}\sim\pi_{\theta}\,(a_{t}\mid\ell,\boldsymbol{o}_{i},a_{<t}).$
@@
For example, OpenVLA first encodes $\boldsymbol{o}_{i}$ through a hybrid vision backbone (DINOv2 [6] w/ SigLIP [51] ) and tokenizes $\ell$ using a pre-trained language model (Llama 2 [43]). These multimodal features are fused in the LLM’s embedding space, which then autoregressively emits $M$ discrete tokens. Each token is finally mapped to an action dimension $a_{t}$ via a de-tokenizer [20].
@@
Diffusion-based Robot Policies. The policy frames action generation as an iterative denoising process [15]. Denote $\boldsymbol{e}_{i}$ the multimodal embeddings fusing $\boldsymbol{o}_{i}$ and $\ell$. Starting from Gaussian noise $\boldsymbol{a}_{i}^{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the reverse diffusion process refines all the $M$ dimensions of the action $\boldsymbol{a}_{i}$ in parallel over $K$ steps through:
@@
$\boldsymbol{a}_{i}^{k-1}=\alpha\left(\boldsymbol{a}_{i}^{k}-\gamma\epsilon_{\theta}\left(\boldsymbol{a}_{i}^{k},\boldsymbol{e}_{i},k\right)\right)+\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ (1)
@@
where $\boldsymbol{a}_{i}^{k}$ is the noisy action at diffusion step $k$, $\epsilon_{\theta}\left(\boldsymbol{a}_{i}^{k},\boldsymbol{e}_{i},k\right)$ is the denoising network, and $\alpha,\gamma,\sigma$ are parameters of a cosine noise schedule [33]. By jointly optimizing the entire $M$ action dimensions of $\boldsymbol{a}_{i}$, diffusion-based robot policies enable fine-grained sampling from the joint probability $\pi_{\theta}(\boldsymbol{a_{i}}|l,\boldsymbol{o}_{i})$ of all action dimensions.
@@ 3.2 Policy Contrastive Decoding (PCD)
The core idea of our proposed PCD method is to eliminate the adverse effect of spurious correlations by redirecting the robot policy’s focus toward object-relevant features during inference. To facilitate the introduction of our PCD, we define spurious correlations as follows. {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt,linecolor=black] Definition 3.1 (Spurious Correlations): Let observations $\boldsymbol{o}$ and language instructions $l$ be generated by a set of $L$ underlying latent factors $\{c_{j}\}_{j=1}^{L}$. These factors are decomposed into task-relevant factors, denoted by the random variable $\boldsymbol{u}$, which exclusively determine the ground-truth expert policy such that $p(\boldsymbol{a}|\boldsymbol{o},l)=p(\boldsymbol{a}|\boldsymbol{u})$, and task-irrelevant factors, denoted by the random variable $\boldsymbol{v}$, which have no causal effect on the expert’s actions $\boldsymbol{a}$. Spurious correlation is then defined as the statistical dependence between the action $\boldsymbol{a}$ and these task-irrelevant factors $\boldsymbol{v}$ observed under the training distribution $p_{\mathrm{train}}(\boldsymbol{a},\boldsymbol{v})$, quantified by the mutual information $I_{\mathrm{train}}(\boldsymbol{a},\boldsymbol{v})$. Spurious correlations, indicated by mutual information $I_{\mathrm{train}}(\boldsymbol{a},\boldsymbol{v})>0$ in the training set, allow a robot policy $\pi_{\theta}$ to infer action $\boldsymbol{a}$ from task-irrelevant factors $\boldsymbol{v}$. These factors act as shortcuts. However, such correlations are unreliable. Under out-of-distribution (OOD) conditions—stemming from environmental shifts or compounding robotic errors—$\boldsymbol{v}$ may change independently of task-relevant factors $\boldsymbol{u}$ and, consequently, $\boldsymbol{a}$, leading to inaccurate action predictions.
@@
A direct approach to mitigate learning from spurious correlations involves filtering out $\boldsymbol{v}$ and utilizing only $\boldsymbol{u}$. Nevertheless, identifying and neutralizing spurious features is challenging due to their significant variability across tasks and scenarios, coupled with interdependencies within the feature space. Drawing inspiration from Vision Contrastive Decoding (VCD) [35, 10], which effectively reduces hallucinations in vision-language models, we introduce Policy Contrastive Decoding (PCD). PCD aims to shift the robot policy’s focus from spurious features $\boldsymbol{v}$ towards task-relevant features $\boldsymbol{u}$ associated with the target object. This is achieved by contrasting model outputs derived from original visual inputs with those from object-masked visual inputs.
@@
Method Overview. Fig. 2 presents an overview of our PCD approach. Suppose we have a pretrained policy $\pi_{\theta}$. Given the current visual observation $\boldsymbol{o}_{i}$ and the language command $\ell$, we first generate two distinct action probability distributions of $\boldsymbol{a}_{i}$: one is $\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ conditioned on the original visual input $\boldsymbol{o}_{i}$, and the other is $\pi_{\theta}(\boldsymbol{a}_{i}|l,\hat{\boldsymbol{o}}_{i})$ conditioned on object-masked counterpart $\hat{\boldsymbol{o}}_{i}$ that removes all task-relevant factors related to target objects. Then, a new contrastive action probability distribution $\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ is obtained by contrasting $\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ and $\pi_{\theta}(\boldsymbol{a}_{i}|l,\hat{\boldsymbol{o}}_{i})$:
$\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})=\frac{1}{C}\cdot\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})\left(\frac{\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})}{\pi_{\theta}(\boldsymbol{a}_{i}|l,\hat{\boldsymbol{o}}_{i})}\right)^{\alpha},$ (2)
@@
where $C$ is a normalization constant and a larger value of $\alpha\geq 0$ indicates a stronger amplification of differences between the two distributions. Particularly, $\alpha=0$ recovers the baseline policy. In such a manner, the obtained $\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ amplifies model predictions on object-relevant features in $\boldsymbol{u}$, thus exhibiting insensitivity to spurious features in $\boldsymbol{v}$, as illustrated in Fig. 2.
@@
To guarantee the flexibility and practical applicability of our PCD method, we need to address the following two questions: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] 1) How can object-masked observations $\hat{\boldsymbol{o}}_{i}$ be automatically generated for each trajectory?
@@
Tracking-to-Mask. We overcome the first question by devising a Tracking-to-Mask (Track2Mask) strategy, which enables precise object masking for sequential visual observations along each trajectory, requiring minimal or no human intervention. Concretely, the first step of Track2Mask is to annotate the task-specified objects in the initial visual observation. Inspired by visual prompting techniques [45, 56], one can annotate the target objects in the initial observation using Point or Box prompts. Besides, we provide the option of leveraging off-the-shelf (open vocabulary) object detection models, such as Grounding DINO [28] for automatic object annotation. Next, the SAM2 [21] model is employed to track and segment the target object across subsequent observations in the trajectory, and the segmented objects are then inpainted to obtain object-masked observations. For more details of Track2Mask, please refer to Appendix A.2.
@@
KDE-based Probabilistic Modeling. We address the second question using KDE-based Probabilistic Modeling (KDE-PM), which approximates action probability distributions for diffusion-based policies through kernel density estimation (KDE) [47]. Specifically, we first use the pretrained diffusion-based policy to sample $N$ candidate action predictions: $\{\boldsymbol{a}_{i}(j)\}_{j=1}^{N}$, where $\boldsymbol{a}_{i}(j)=[{a}_{1}(j),...,{a}_{M}(j)]\sim\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$. With the assumption that all action dimensions are independent, i.e., $\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})=\prod\limits_{t=1}^{M}\pi_{\theta}(a_{t}|l,\boldsymbol{o}_{i})$, we calculate an approximated probability distribution for each action dimension of the action $\boldsymbol{a}_{i}$ via KDE:
$\pi_{\theta}(a_{t}|l,\boldsymbol{o}_{i})\approx\frac{1}{C^{\prime}}\sum_{j=1}^{N}\mathcal{K}\left(\frac{a_{t}-a_{t}(j)}{b}\right),\,\,t=1,...,M,$ (3)

### v3

@@ Abstract
Robotic foundation models, or generalist robot policies, hold immense potential to enable flexible, general-purpose and dexterous robotic systems. Despite their advancements, our empirical experiments reveal that existing robot policies are prone to learning spurious correlations from pre-training trajectories, adversely affecting their generalization capabilities beyond the training data. To tackle this, we propose a novel Policy Contrastive Decoding (PCD) approach, which redirects the robot policy’s focus toward object-relevant visual clues by contrasting action probability distributions derived from original and object-masked visual inputs. As a training-free method, our PCD can be used as a plugin to improve different types of robot policies without needing to finetune or access model weights. We conduct extensive experiments on top of three open-source robot policies, including the autoregressive policy OpenVLA and the diffusion-based policies Octo and $\pi_{0}$. The obtained results in both simulation and real-world environments prove PCD’s flexibility and effectiveness, e.g., PCD enhances the state-of-the-art policy $\pi_{0}$ by 8.9% in the simulation environment and by 108% in the real-world environment.
@@ 1 Introduction
Recent efforts in developing flexible, general-purpose, and dexterous robotic systems have largely focused on robotic foundation models, or generalist robot policies. The goal of such policies is to empower users to instruct robots to perform arbitrary tasks, enabling autonomous task execution with minimal human supervision. Specifically, the policy takes as input a visual observation of the robot’s state combined with a language instruction that defines the task, and outputs a robot action, such as end-effector displacement. Ongoing advancements in scaling real-world robotic data corpora have facilitated the development of numerous robot policies (Brohan et al., 2022; Kim et al., 2024; M. et al., 2024; Pertsch et al., 2025), which have demonstrated exceptional effectiveness in controlling various robots across diverse environments and acquiring a wide range of manipulation skills.
[…]
@@
In this work, we present a Policy Contrastive Decoding (PCD) approach to answer the question. The core idea of our PCD approach is to shift the robot policy’s attention from spurious features to object-relevant ones by contrasting action probability distributions derived from original and object-masked observation images during inference. To ensure both effectiveness and flexibility of the approach, a Tracking-to-Mask (Track2Mask) strategy is introduced, which first annotates target-specific objects in the initial visual observation using human expertise (e.g., Point and Box prompts) or off-the-shelf object detection models (e.g., Grounding DINO (Liu et al., 2024b)), then employs the SAM2 (Kirillov et al., 2023) model to track these objects across subsequent observations in the trajectory. This enables precise object masking with minimal or no human intervention. Moreover, unlike autoregressive policies, diffusion-based policies cannot directly generate action probability distributions to feed into our PCD. To cope with this limitation, we introduce a KDE-based Probabilistic Modeling (KDE-PM) scheme, which calculates action probability distributions for diffusion-based policies using the kernel density estimation (KDE) (Wkeglarczyk, 2018) technique. This enables our PCD approach to be compatible with both autoregressive and diffusion-based policies.
@@
Figure 1: Robot policies tend to spuriously correlate task-irrelevant features with actions, compromising their ability to generalize to unseen scenarios. As observed, changing the light position from (a) to (b) and the drawer handle position from (a) to (c) results in 36% and 32% drops in the performance of the baseline policy OpenVLA (Kim et al., 2024), respectively. (d) Attention map. More results are in Section 4.4 and Appendix A.
@@
To the best of our knowledge, this is the first work to introduce a training-free approach for resolving the spurious correlation issue in robot policies. We conduct extensive evaluations in both simulation and real-world environments, across a total of 15 diverse tasks, and on top of three open-source robot policies—including the autoregressive policy OpenVLA (Kim et al., 2024) and the diffusion-based policies Octo (M. et al., 2024) and $\pi_{0}$ (Black et al., 2024). The achieved results demonstrate the strong flexibility and effectiveness of our PCD approach. On simulation benchmarks, PCD enhances the performance of Octo, OpenVLA and $\pi_{0}$ by 29.7%, 50.6% and 8.9%, respectively. On real-world manipulation tasks, PCD outperforms the state-of-the-art robot policy $\pi_{0}$ by 108%.
@@
Contributions. Our contributions in this work are threefold. 1) We propose PCD, a simple, training-free, and easy-to-implement scheme to tackle the spurious correlation issue in robot policies. 2) PCD can be used as a plugin to enhance both autoregressive and diffusion-based policies, without requiring fine-tuning or access to pre-trained model weights. 3) Extensive experiments demonstrate PCD’s effectiveness and flexibility, consistently improving three state-of-the-art policies across 15 diverse tasks.
@@ 2 Related Work
Robotic Foundation Models. The rapid development of robotic foundation models or generalist robot policies has been significantly inspired by the success of large visual and language models (Karamcheti et al., 2024a; Zhai et al., 2023a; Touvron et al., 2023a). Robot policies can be categorized based on action prediction strategies into autoregressive models and diffusion-based policies. Autoregressive policies sequentially generate action tokens based on previous outputs (Brohan et al., 2022; 2023; Zawalski et al., 2024; Belkhale et al., 2024; Zhang et al., 2024). OpenVLA (Kim et al., 2024) equips with a Llama 2 (Touvron et al., 2023a) language model alongside a visual encoder that merges pretrained features from SigLIP (Zhai et al., 2023a) and DINO-v2 (Oquab et al., 2024). Diffusion-based policies simultaneously generate entire action dimensions by performing diffusion processes on noise (M. et al., 2024; Liu et al., 2024a). Octo (Team et al., 2024) learns a lightweight policy on the OXE dataset (Collaboration et al., 2024), enabling rapid adaptation to unknown tasks and scenarios on standard consumer-grade GPUs. $\pi_{0}$ (Black et al., 2024) exhibits flexible control over different robot types by leveraging pre-trained general knowledge from vision-language models. Although most robot policies are trained on extensive robotic behavior datasets, recent studies have shown that slight variations in deployment scenarios can significantly affect their generalization performance on downstream tasks in real-world environments (Firoozi et al., 2023; Gao et al., 2024).
@@
Contrastive Decoding. Recent studies in large vision-language models (VLMs) have highlighted the challenge of hallucination (Gunjal et al., 2024; Rawte et al., 2023; Liu et al., 2023b), where models generate inaccurate or misleading outputs not grounded in the input data. Contrastive Decoding (CD) (Favero et al., 2024; Chen et al., 2024; Leng et al., 2024) suppresses hallucinations by comparing output distributions from original and distorted input without altering the model’s weights. Instruction Contrastive Decoding (ICD) (Wang et al., 2024) is used to combat hallucinations in multimodal tasks by modulating the confidence in multimodal alignment of the model’s visual and textual inputs, enabling it to distinguish between hallucinated and relevant tokens. Visual Contrastive Decoding (VCD) (Park et al., 2025; Favero et al., 2024) aims to reduce object hallucinations by comparing outputs from original and distorted visual inputs. This approach provides a computationally efficient solution as it requires neither additional training nor external pre-trained models. Our proposed Policy Contrastive Decoding (PCD) approach in this work is mainly inspired by the success of VCD schemes. However, unlike those schemes that contrast output distributions from original and noise-distorted (Park et al., 2025) or image-removed (Favero et al., 2024) inputs to counteract LVLMs’ overreliance on language priors, our PCD contrasts action probability distributions derived from original and object-masked inputs, thereby redirecting the robot policy’s focus to object-relevant visual features and enhancing their generalization across different scenarios.
@@ 3.1 Preliminaries
This work focuses on robot manipulation tasks where the objective is to leverage a learned language-conditioned policy $\pi_{\theta}(\bm{a}_{i}\mid\bm{o}_{i},\ell)$ to predict an $M$-dimensional action $\bm{a}_{i}=[a_{1},...,a_{M}]\in\mathbb{R}^{M}$ conditioned on the current visual observation $\bm{o}_{i}$ and the language instruction $\ell$ of the task. In this part, we formalize two mainstream paradigms for action prediction: autoregressive action prediction (e.g., OpenVLA (Kim et al., 2024)) and diffusion-based action prediction (e.g., Octo (M. et al., 2024) and $\pi_{0}$ (Black et al., 2024)).
@@
Figure 2: Overview of our Proposed Policy Contrastive Decoding (PCD) approach. PCD serves as a plugin to redirect the robot policy’s focus toward object-relevant visual cues by contrasting action probability distributions derived from original observations p{p} and object-masked observations $\hat{p}$. For illustrative purposes, we visualize the predictions only in the $\Delta x$ and $\Delta y$ dimensions of the robot action space $\Delta x$, $\Delta y$, $\Delta z$, rotx, roty, rotz, gripper].
@@
Autoregressive Robot Policies. The policy decomposes the action prediction task into sequential token generation. Each action dimension $a_{t}$ of $\bm{a}_{i}$ is sampled autoregressively from the probability distribution, conditioned on the visual observation $\bm{o}_{i}$, the language instruction $\ell$, as well as the previously predicted dimensions $a_{<t}=\{{a}_{t}\}_{t=1}^{t-1}$:
$a_{t}\sim\pi_{\theta}\,(a_{t}\mid\ell,\bm{o}_{i},a_{<t}).$
@@
For example, OpenVLA first encodes $\bm{o}_{i}$ through a hybrid vision backbone (DINOv2 (Caron et al., 2023) w/ SigLIP (Zhai et al., 2023b) ) and tokenizes $\ell$ using a pre-trained language model (Llama 2 (Touvron et al., 2023b)). These multimodal features are fused in the LLM’s embedding space, which then autoregressively emits $M$ discrete tokens. Each token is finally mapped to an action dimension $a_{t}$ via a de-tokenizer (Kim et al., 2024).
@@
Diffusion-based Robot Policies. The policy frames action generation as an iterative denoising process (Ho et al., 2020). Denote $\bm{e}_{i}$ the multimodal embeddings fusing $\bm{o}_{i}$ and $\ell$. Starting from Gaussian noise $\bm{a}_{i}^{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the reverse diffusion process refines all the $M$ dimensions of the action $\bm{a}_{i}$ in parallel over $K$ steps through:
@@
$\bm{a}_{i}^{k-1}=\alpha\left(\bm{a}_{i}^{k}-\gamma\epsilon_{\theta}\left(\bm{a}_{i}^{k},\bm{e}_{i},k\right)\right)+\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ (1)
@@
where $\bm{a}_{i}^{k}$ is the noisy action at diffusion step $k$, $\epsilon_{\theta}\left(\bm{a}_{i}^{k},\bm{e}_{i},k\right)$ is the denoising network, and $\alpha,\gamma,\sigma$ are parameters of a cosine noise schedule (Nichol and Dhariwal, 2021). By jointly optimizing the entire $M$ action dimensions of $\bm{a}_{i}$, diffusion-based robot policies enable fine-grained sampling from the joint probability $\pi_{\theta}(\bm{a_{i}}|l,\bm{o}_{i})$ of all action dimensions.
@@ 3.2 Policy Contrastive Decoding (PCD)
The core idea of our proposed PCD method is to eliminate the adverse effect of spurious correlations by redirecting the robot policy’s focus toward object-relevant features during inference. To facilitate the introduction of our PCD, we define spurious correlations as follows. Definition 3.1 (Spurious Correlations): Let observations $\bm{o}$ and language instructions $l$ be generated by a set of $L$ underlying latent factors $\{c_{j}\}_{j=1}^{L}$. These factors are decomposed into task-relevant factors, denoted by the random variable $\bm{u}$, which exclusively determine the ground-truth expert policy such that $p(\bm{a}|\bm{o},l)=p(\bm{a}|\bm{u})$, and task-irrelevant factors, denoted by the random variable $\bm{v}$, which have no causal effect on the expert’s actions $\bm{a}$. Spurious correlation is then defined as the statistical dependence between the action $\bm{a}$ and these task-irrelevant factors $\bm{v}$ observed under the training distribution $p_{\mathrm{train}}(\bm{a},\bm{v})$, quantified by the mutual information $I_{\mathrm{train}}(\bm{a},\bm{v})$.
[…]
@@
A direct approach to mitigate learning from spurious correlations involves filtering out $\bm{v}$ and utilizing only $\bm{u}$. Nevertheless, identifying and neutralizing spurious features is challenging due to their significant variability across tasks and scenarios, coupled with interdependencies within the feature space. Drawing inspiration from Vision Contrastive Decoding (VCD) (Park et al., 2025; Favero et al., 2024), which effectively reduces hallucinations in vision-language models, we introduce Policy Contrastive Decoding (PCD). PCD aims to shift the robot policy’s focus from spurious features $\bm{v}$ towards task-relevant features $\bm{u}$ associated with the target object. This is achieved by contrasting model outputs derived from original visual inputs with those from object-masked visual inputs.
@@
Method Overview. Fig. 2 presents an overview of our PCD approach. Suppose we have a pretrained policy $\pi_{\theta}$. Given the current visual observation $\bm{o}_{i}$ and the language command $\ell$, we first generate two distinct action probability distributions of $\bm{a}_{i}$: one is $\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})$ conditioned on the original visual input $\bm{o}_{i}$, and the other is $\pi_{\theta}(\bm{a}_{i}|l,\hat{\bm{o}}_{i})$ conditioned on object-masked counterpart $\hat{\bm{o}}_{i}$ that removes all task-relevant factors related to target objects. Then, a new contrastive action probability distribution $\pi_{\theta}^{*}(\bm{a}_{i}|l,\bm{o}_{i})$ is obtained by contrasting $\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})$ and $\pi_{\theta}(\bm{a}_{i}|l,\hat{\bm{o}}_{i})$:
$\pi_{\theta}^{*}(\bm{a}_{i}|l,\bm{o}_{i})=\frac{1}{C}\cdot\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})\left(\frac{\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})}{\pi_{\theta}(\bm{a}_{i}|l,\hat{\bm{o}}_{i})}\right)^{\alpha},$ (2)
@@
where $C$ is a normalization constant and a larger value of $\alpha\geq 0$ indicates a stronger amplification of differences between the two distributions. Particularly, $\alpha=0$ recovers the baseline policy. In such a manner, the obtained $\pi_{\theta}^{*}(\bm{a}_{i}|l,\bm{o}_{i})$ amplifies model predictions on object-relevant features in $\bm{u}$, thus exhibiting insensitivity to spurious features in $\bm{v}$, as illustrated in Fig. 2.
@@
To guarantee the flexibility and practical applicability of our PCD method, we need to address the following two questions: 1) How can object-masked observations $\hat{\bm{o}}_{i}$ be automatically generated for each trajectory?
@@
Tracking-to-Mask. We overcome the first question by devising a Tracking-to-Mask (Track2Mask) strategy, which enables precise object masking for sequential visual observations along each trajectory, requiring minimal or no human intervention. Concretely, the first step of Track2Mask is to annotate the task-specified objects in the initial visual observation. Inspired by visual prompting techniques (Wan et al., 2024), one can annotate the target objects in the initial observation using Point or Box prompts. Besides, we provide the option of leveraging off-the-shelf (open vocabulary) object detection models, such as Grounding DINO (Liu et al., 2024b) for automatic object annotation. Next, the SAM2 (Kirillov et al., 2023) model is used to track and segment the target object across subsequent observations in the trajectory, and the segmented objects are then inpainted to obtain object-masked observations. For more details of Track2Mask, please refer to Appendix B.
@@
KDE-based Probabilistic Modeling. We address the second question using KDE-based Probabilistic Modeling (KDE-PM), which approximates action probability distributions for diffusion-based policies through kernel density estimation (KDE) (Wkeglarczyk, 2018). Specifically, we first use the pretrained diffusion-based policy to sample NN candidate action predictions: $\{\bm{a}_{i}(j)\}_{j=1}^{N}$, where $\bm{a}_{i}(j)=[{a}_{1}(j),...,{a}_{M}(j)]\sim\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})$. With the assumption that all action dimensions are independent, i.e., $\pi_{\theta}(\bm{a}_{i}|l,\bm{o}_{i})=\prod\limits_{t=1}^{M}\pi_{\theta}(a_{t}|l,\bm{o}_{i})$, we calculate an approximated probability distribution for each action dimension of the action $\bm{a}_{i}$ via KDE:
$\pi_{\theta}(a_{t}|l,\bm{o}_{i})\approx\frac{1}{C^{\prime}}\sum_{j=1}^{N}\mathcal{K}\left(\frac{a_{t}-a_{t}(j)}{b}\right),\,\,t=1,...,M,$ (3)